# LAMB: LLM-based Audio Captioning with Modality Gap Bridging via Cauchy-Schwarz Divergence

### v1

@@ 1 Introduction
Automated Audio Captioning (AAC) is a multimodal task that aims to generate corresponding descriptions for a given audio content [36]. Capturing acoustic semantic information and aligning it with textual features are key challenges in AAC, as these processes enable the model to generate semantically consistent and contextually appropriate descriptions.
Prior works [19, 33, 18, 31, 23, 4, 17, 14, 32, 6] typically follow a two-stage framework, using an audio encoder [20, 3, 12] and a text decoder [9, 30, 21]. More recently, with the emergence of large language models (LLMs), studies [33, 23, 18, 31, 4] have employed LLMs [5, 34, 7] as a text decoder to leverage their contextual reasoning.
However, because LLMs are trained on large-scale text corpora, their reasoning is most effectively exploited when the conditioning inputs (e.g. audio modality) are represented in the LLM text embedding space [25]. Existing methods [33, 23, 4] map audio features into this space, typically via linear projection or Q-Former [22], without an explicit objective that bridges the audio–text modality gap. As a result, a modality misalignment can remain between audio representations and the LLM embedding space, limiting the ability of the decoder that interprets audio semantics.
@@ 2.1 Two-Stream Adapter
$\mathrm{MHA}(\mathbf{Q}_{s},\mathbf{h}_{a})=\sigma\left(\frac{\textbf{Q}_{s}W_{Q}\cdot(\mathbf{h}_{a}W_{k})^{\top}}{\sqrt{d_{H}}}\right)\mathbf{h}_{a}W_{v},$
$\hat{\mathbf{h}}_{s}=\mathrm{LN}(\mathrm{MHA}(\mathbf{Q}_{s},\mathbf{h}_{a})+\mathbf{Q}_{s})),$
@@
Fusion. The outputs of the two parallel modules $\hat{h}_{s}$ and $\hat{h}_{t}$ are concatenated along the token axis, and then processed with $N_{g}$ global queries $\mathbf{Q}_{g}$ via a cross-attention layer with residual and LayerNorm. The resulting representations are projected into the $D_{llm}$-dimensional LLM text embedding space via $W_{p}\in\mathbb{R}^{D\times D_{llm}}$, yielding $\mathbf{z}_{a}\in\mathbb{R}^{N_{g}\times D_{llm}}$:
@@ 3.1.2 Implementation Details and Metrics
We adopt the consistent ensemble distillation [10] model as the audio encoder and LLaMA 2 (7B) [34] as the text decoder, fine-tuned with LoRA [15]. Query numbers are set to $N_{s}$=8, $N_{t}$=8, and $N_{g}$=32. All experiments used the AdamW with weight decay 1e-6, trained for 30 epochs (2 warmup) under a cosine annealing schedule. Pre-training employed a learning rate of 5e-5 with batch size 32, and fine-tuning used 3e-6 with batch size 8. For evaluation, we adopt common AAC metrics, including METEOR [2], CIDEr [35], SPICE [1], SPIDEr [24], SPIDEr-FL [38], and FENSE [38].
@@ 3.2 Main Results
Tab. LABEL:tab:main_tab summarizes the performance of LAMB on the AAC benchmarks. LAMB surpasses prior work on all evaluation metrics, establishing clear state-of-the-art results on AudioCaps. On AudioCaps, it achieves notable gains, particularly on CIDEr, SPIDEr, and SPIDEr-FL, while also improving METEOR and SPICE, reflecting stronger semantic and structural fidelity. On Clotho, which is more challenging due to diverse captions, LAMB shows competitive performance with improvements on several metrics, demonstrating robustness to annotation variability. Notably, while several prior works rely on additional pre-training data, our model attains comparable or better results across the dataset.
@@ 3.4 Qualitative Results
In this work, we introduced LAMB, an LLM-based audio captioning framework that applies Cauchy–Schwarz (CS) divergence to bridge the audio–text modality gap for the first time. LAMB integrates a Cross-Modal Aligner for fine-grained alignment and a Two-Stream Adapter to enrich audio representations, while the Token Guide steers decoder logits using guide scores in the LLM text embedding space. Experimental results show that LAMB outperforms prior work on AudioCaps and achieves competitive performance on Clotho, indicating that bridging the modality gap enables more effective use of LLM reasoning for audio understanding.

### v2

@@ 1 Introduction
Automated Audio Captioning (AAC) is a multimodal task that aims to generate corresponding descriptions for given audio content [37]. Capturing acoustic semantic information and aligning it with textual features are key challenges in AAC, as these processes enable the model to generate semantically consistent and contextually appropriate descriptions.
Prior works [19, 34, 18, 32, 23, 4, 17, 14, 33, 6] typically follow a two-stage framework, using an audio encoder [20, 3, 12] and a text decoder [9, 31, 21]. More recently, with the emergence of large language models (LLMs), studies [34, 23, 18, 32, 4, 27] have employed LLMs [5, 35, 7] as a text decoder to leverage their contextual reasoning.
However, because LLMs are trained on large-scale text corpora, their reasoning is most effectively exploited when the conditioning inputs (e.g., audio modality) are represented in the LLM text embedding space [25]. Existing methods [34, 23, 4] map audio features into this space, typically via linear projection or Q-Former [22], without an explicit objective that bridges the audio–text modality gap. As a result, a modality misalignment can remain between audio representations and the LLM embedding space, limiting the ability of the decoder that interprets audio semantics.
@@ 2.1 Two-Stream Adapter
$\mathrm{MHA}(\mathbf{Q}_{s},\mathbf{h}_{a})=\sigma\left(\frac{\mathbf{Q}_{s}W_{Q}\cdot(\mathbf{h}_{a}W_{K})^{\top}}{\sqrt{d_{H}}}\right)\mathbf{h}_{a}W_{V},$
$\hat{\mathbf{h}}_{s}=\mathrm{LN}(\mathrm{MHA}(\mathbf{Q}_{s},\mathbf{h}_{a})+\mathbf{Q}_{s}),$
@@
Fusion. The outputs of the two parallel modules $\hat{\mathbf{h}}_{s}$ and $\hat{\mathbf{h}}_{t}$ are concatenated along the token axis, and then processed with $N_{g}$ global queries $\mathbf{Q}_{g}$ via a cross-attention layer with residual and LayerNorm. The resulting representations are projected into the $D_{llm}$-dimensional LLM text embedding space via $W_{p}\in\mathbb{R}^{D\times D_{llm}}$, yielding $\mathbf{z}_{a}\in\mathbb{R}^{N_{g}\times D_{llm}}$:
@@ 3.1.2 Implementation Details and Metrics
We adopt the consistent ensemble distillation [10] model as the audio encoder and LLaMA 2 (7B) [35] as the text decoder, fine-tuned with LoRA [15]. Query numbers are set to $N_{s}$=8, $N_{t}$=8, and $N_{g}$=32. All experiments used AdamW with weight decay 1e-6, trained for 30 epochs (2 warmup epochs) under a cosine annealing schedule. Pre-training employed a learning rate of 5e-5 with batch size 32, and fine-tuning used 3e-6 with batch size 8. For evaluation, we adopt common AAC metrics, including METEOR [2], CIDEr [36], SPICE [1], SPIDEr [24], SPIDEr-FL [39], and FENSE [39].
@@ 3.2 Main Results
Tab. LABEL:tab:main_tab summarizes the performance of LAMB on the AAC benchmarks. LAMB surpasses prior work on all evaluation metrics, establishing clear state-of-the-art results on AudioCaps. On AudioCaps, it achieves notable gains, particularly on CIDEr, SPIDEr, and SPIDEr-FL, while also improving METEOR and SPICE, reflecting stronger semantic and structural fidelity. On Clotho, which is more challenging due to diverse captions, LAMB shows competitive performance with improvements on several metrics, demonstrating robustness to annotation variability. Notably, while several prior works rely on additional pre-training data, our model attains comparable or better results across the datasets.
@@ 4 Conclusion
In this work, we introduced LAMB, an LLM-based audio captioning framework that, for the first time, applies Cauchy-Schwarz divergence to bridge the audio-text modality gap in audio captioning. LAMB integrates a Cross-Modal Aligner for fine-grained alignment and a Two-Stream Adapter to enrich audio representations, while the Token Guide steers decoder logits using Guide Scores in the LLM text embedding space. Experimental results show that LAMB outperforms prior work on AudioCaps and achieves competitive performance on Clotho, indicating that bridging the modality gap enables effective use of LLM reasoning in audio captioning.
5 Acknowledgement
This work was supported by Institute of Information &\& communications Technology Planning &\& Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2025-02215122, Development and Demonstration of Lightweight AI Model for Smart Homes).